A heavy mass on a frictionless ramp at an angle of 30 degrees drags a 2 kg mass upward using a pulley. if the tension in the rope is 30 N, what is the acceleration of the 2 kg mass?
1
Expert's answer
2010-11-19T08:39:43-0500
Let's write II Newton's law for the mass:
ma = T - mg sin(α);
a = T/m - g sin(α) = 30/2 - 10 sin(30) = 10 m/s2